! Required information The Foundational 15 (Algo) (LO5-1, LO5-3, LO5-4, LO5-5, LO5-6, LO5-7, LO5-8] [The following informatioRequired: 1. What is the contribution margin per unit? (Round your answer to 2 decimal places.) Contribution margin per unit2. What is the contribution margin ratio? Contribution margin ratio %3. What is the variable expense ratio? Variable expense ratio %4. If sales increase to 1,001 units, what would be the increase in net operating income? (Round your answer to 2 decimal plac5. If sales decline to 900 units, what would be the net operating income? Net operating income
Answer #1

Sales Per Unit = (Sales/units) = $50,000/1,000 = $50. Variable Cost per unit = (Variable cost/Unit) = $27,500/1,000 = $27.50.

1) Contribution Margin per Unit = Sales per unit - Variable cost per Unit => $50 - $27.50 = $22.50.

2)Contribution Margin ratio = (Contribution margin/Sales) => ($22.50/$50)*100 = 45%.

3) Variable Cost ratio = 100 - Contribution margin ratio => (100 - 45%) = 55%.

4)Increase in net Operating income.

Net Income at 1,001 units = (Sales - Variable cost per unit) - Fixed cost => ($50 - $27.50)*1,001 - $14,850 =>$7,672.50.

Net Operating Income at 1000 units = $7,650.

Net Increase in net Operating Income = $7,672.50 - $7,650 = $22.50.

5) Net Operating Income at 900 Units.

=> Sales Per unit - Variable Cost per unit - Fixed Cost. =>[($50 - $27.50) * 900] - $14,850 = $5,400.

= $5,400.

If you have any doubts please comment on the answer.